# Speedups in nonequilibrium thermal relaxation: Mpemba and related effects

### v1

@@ 2.1.7 Mpemba effect in boundary and weak coupling setups
To model bulk transitions, we use rates that decay exponentially as $2^{-d_{i\!j}}$, where $d_{i\!j}=\sum_{s}\delta_{\sigma_{s}^{i},-\sigma_{s}^{j}}$ is the Hamming distance [81] that counts the number of spins that must be flipped between the two configurations, as suggested by decimation-like procedures operated on Markov jump processes [78]. We, therefore formalize bulk transitions between two states $i\neq j$ as
The full transition matrix for the model is finally built as a linear combination of the two rate matrices, as in Eq. (43). With this construction, let us consider the persistence of the Mpemba effect in a boundary-coupling setup. In Fig. 11b, we plot the Mpemba phase diagram for $N=10$ (implying $M=1024$ microstates) as a function of bath temperature $\beta_{b}$ and external magnetic field $H$. The resemblance with that of the model studied in [5] is striking, especially considering the profound dissimilarities between the two models: the all-to-all connectivity of a mean-field description makes it relevant for high-dimensional setups, while the model presented here addresses a one-dimensional spin chain. Additionally, in this model, we imposed a partial coupling with the thermal bath, while in the standard mean-field Ising antiferromagnet, every spin exchanges heat with the thermal bath. The similarity is not limited only to the examples where the Mpemba effect can be observed but extends to the many specific flavors of the Mpemba effect; see Fig. 12 for a detailed analysis. In general, we observe both direct/inverse, weak/strong and single/multiple effects, which are encoded by the different colors in the figure: here, boundary coupling merely shifts phase-diagram borders, and all flavors of the Mpemba effect can be observed. This hints that the antiferromagnetic character lies at the origin of the Mpemba effect, as it is the only feature shared among these very different models. The horizontal line at $H=2$ is determined by the phase transition that characterizes the system at $T_{b}=0$: for $0<H\leq 2$ (alternating spins ground state), only single Mpemba effects can be observed, while for $H>2$ (aligned spins ground state) double effects occur. For $0<H\leq 2$, a strong inverse Mpemba effect effect exists for $T_{b}\geq T_{b}^{*}=1/\beta^{*}$ (highlighted with a red arrow). The reason that no Mpemba effect exists for $T<T_{b}^{*}$, for small external magnetic fields, traces back to a phase transition of the relaxation dynamics [82]; see the discussion of applications in Sec. 8.
Additional analysis that demonstrates the Mpemba effect in this setup can be obtained by projecting the relaxation paths in the (high-dimensional) probability space onto two order parameters: the average $\langle m\rangle$ and staggered magnetization $\langle m_{s}\rangle$ [65]. This projection method, explained in Sec. 7.1.3, allows Monte-Carlo simulation in large systems where a direct calculation of $a_{2}$ is impractical. We can nonetheless compare the relaxation trajectory with the equilibrium distributions at different temperatures; see Sec. 7.1.3 for an example in a 2D Ising antiferromagnet.
Figure 13 illustrates different relaxation paths in a weakly coupled chain of $N=50$ spins in comparison to the equilibrium line. Initial temperatures are chosen before, at and after the zero of $a_{2}$ at $T_{0}\sim 0.6<T_{b}$. The hot (red) and cold (blue) relaxation paths approach the bath equilibrium at $T_{b}=5$ along the same direction, which is the projection of $\vec{v}_{2}$ into the observable space, but with opposite direction. This corresponds to a sign-change in $a_{2}$, which indeed vanishes at the blue trajectory implying nonmonotonicity of $a_{2}(T_{0})$ (inset of Fig. 13) and consequently a Mpemba effect in the system.
Figure 13: Relaxation paths for a weakly coupled system ($N=50$) on the $m\right>$-$m_{s}\right>$ plane ($H=\sqrt{2}$). Initial temperatures are chosen before (blue), at (green) and after (red) the zero of $a_{2}(T_{0})$ as shown in the inset. Red and green paths approach the equilibrium locus along the same vector but from opposite directions, indicating the existence of a Mpemba effect. The figure is adapted from [70].
@@
Figure 14: (a) Potential $U(x)$ of the colloidal system used in [64]. (b) Examples of Boltzmann distributions at cold and hot bath temperatures for which the direct (cooling) and inverse (heating) Mpemba effect exist. (c) Mpemba phase diagram as a function of the bath temperature $T_{b}$ and thermal diffusivity $\kappa$. The figures are taken from [70].
@@
$\partial_{t}p(x,t)=-\mu\partial_{x}\Big{[}\big{(}\partial_{x}U(x)\big{)}p(x,t)\Big{]}+\mu\partial_{x}^{2}\Big{(}T(x,t)p(x,t)\Big{)},$ (47)
where $U(x)$ is the potential and $\mu$ is the mobility of the Brownian particle. Although the Fokker-Plank operator is time-dependent, the long-time limit $T(x)\to T_{b}$ implies $\mathcal{L}(t)\to\mathcal{L}_{T_{b}}$ and ensures convergence to the Boltzmann equilibrium $\pi(x,T_{b})\propto e^{-U(x)/k_{B}T_{b}}$. The eigenfunctions of $\mathcal{L}_{T_{b}}$, solving $\mathcal{L}_{T_{b}}v_{i}(x,T_{b})=\lambda_{i}v_{i}(x,T_{b})$ with $0=\lambda_{1}>\lambda_{2}\geq\lambda_{3}\geq\dots,$ form a complete basis; as a result,
@@ 2.1.8 Metastability
The maximum amount of work that can be extracted from a nonequilibrium isothermal protocol is the equality of Eq. (50), and the authors refer to it as the extractable work,
@@
At the end of the first stage, the particle has quickly descended into a well. The local equilibrium probability density, $\pi_{{\rm leq}}(x,T,T_{b})$, locally looks like the (global) equilibrium density $\pi(x,{T_{b}})$ yet with altered statistical weight in the left and right wells,
@@
The shape of the potential plays an important role in determining whether there is a Mpemba effect. In this discussion, we focused on cases with metastability in the system, but other aspects of the potential’s geometry are also significant. For example, in [3] a proposal for a one-dimensional potential that, under certain initial conditions and with overdamped diffusion dynamics, exhibits the Mpemba effect. Unlike the above discussion, this potential features (in addition to a ground state and a metastable state) and a wide, shallow well (see Fig. 6). This energy landscape facilitates both a “fast” and a “slow” direction for relaxation. The relaxation from the metastable state is “sluggish” due to the depth of that well. Starting at a high temperature is advantageous because the majority of the statistical weight resides in the shallow parts of the potential, allowing for a “fast” relaxation to the ground state.
The presence of metastability significantly affects the type and occurrence of the Mpemba effect, as illustrated in both the spin-glass example in Section 7.1 and the Random Energy Model discussed in Section 5. However, metastability is not a prerequisite for observing the Mpemba effect. Walker and Vucelja in [91] examined an overdamped particle diffusing within a piecewise-continuous potential characterized by energetically flat regions. In these flat regions, the particle can diffuse freely but has a high entropy, making it time-consuming to “escape” from them. This piecewise-continuous potential can be coarse-grained into three states, where the energy levels correspond to different free-energy levels, leading to a form of ”hidden” metastability. A generalization of the above results to piecewise linear potential was done by Biswas, Rajesh, and Pall in [92] as well as in [93]. They investigated analytically a piecewise-linear double potential well [92] (as well as the single well, with and without activity, in [93]) and concluded that neither metastability nor the asymmetry of the potential is necessary or sufficient for the Mpemba effect to occur. A similar conclusion appeared in Kumar’s thesis [94], where it was numerically demonstrated that the Mpemba effect exists in a single-well potential without metastability.
@@
Figure 15: Probability density dynamics in the space of projection coefficients, for the double-well potential described in [6, 75]. (a) Two-dimensional projection of the dynamics onto coefficients $a_{2}$ and $a_{3}$. Red line, labeled $G$, denotes the locus of global equilibrium probability densities (equilibria); green line, labeled $G_{\rm leq}$, denotes the locus of local equilibria. Gray lines denote individual trajectories starting from initial temperatures $T_{0}=\{1000,100,\ldots\}$, labeled by black markers, to the global equilibrium at $T_{b}=1$ (large hollow marker with central dot). The time evolution of the trajectories is indicated by a color gradient from dark to light gray. Time along the local equilibrium locus (green line) is proportional to the relaxation time, as the initial stage is fast. The nonmonotonicity of initial temperatures with respect to length along $G_{\rm leq}$ indicates a Mpemba effect. (b) Three-dimensional projection onto the space of $a_{2}$-$a_{3}$-$a_{4}$ coefficients shows that the red and green curves intersect only at $T_{0}=T_{b}$. Both panels are based on data from experiments [95]. Figure is reproduced from [75].
@@ 2.1.9 The role of dynamics
In some systems, e.g., the 1D anti-ferromagnetic system discussed above, the existence of some type of Mpemba effect seems natural from the shape of the equilibrium line; see, for example, Fig. 13. It is important to stress, however, that this is not the case: the existence or non-existence of these effects is a consequence of the dynamics and cannot be deduced from the equilibrium properties of the system alone. This is because the effect is based on the projection of the equilibrium distribution on $\vec{v}_{2}$, which is a property of the dynamics, and is not an equilibrium property. In contrast to the equilibrium properties, the exact dynamics are non-universal and affected by many specific issues, such as the precise type of coupling between the system and the bath.
One might imagine that if the equilibrium line is sufficiently complex, the equilibrium projections on $\vec{v}_{2}$ could be nonmonotonic, regardless of the direction of $\vec{v}_{2}$. However, this is never the case: for any system that obeys detailed balance, there is at least one dynamical law in which no Mpemba effect occurs. In fact, a fully connected graph with Metropolis dynamics does not exhibit any type of Mpemba effect, regardless of the equilibrium distribution [96]. Interestingly, in some cases, removing a single link, which corresponds to introducing an infinite barrier, or two links in a fully connected graph with Metropolis dynamics can be enough to induce the Mpemba effect [97]. Similarly, in [98], it was shown that in chemical reactions with identical initial conditions and the same equilibrium target distribution, the existence of the Mpemba effect depends on the precise details of the dynamics. Altering the dynamics, in some cases, one can make the effect appear or disappear. Such implications of dynamics in the context of the rates of chemical reactions are discussed in Sec. 8.1.4. Additionally, applications related to optimal transport and engine efficiency can be found in Sec. 8.1.2.
@@ 2.2 Quantum systems
Until recently, the Mpemba effect was studied exclusively in the context of classical systems. Indeed, a 2018 review of “Quantum quench dynamics” [99] has no mention of a Mpemba or even a Mpemba-related effect. Yet in the last five years, the situation has changed dramatically, and this review includes over fifty references that relate to the Mpemba effect or analogs seen in quantum systems. It is largely believed that none of the quantum effects are the origin of the classical effects. Instead, these are quantum effects that share properties that are similar to classical effects. Much as in classical systems, several different quantum effects have been considered, and some were observed experimentally. Below, we give a brief conceptual overview of this fast-developing field. We discuss experiments from 2024 in Sec. 6.
@@ 2.2.1 Theoretical framework: Markovian open quantum systems
Since the dynamics generated by $\mathcal{L}$ is completely positive, all the eigenvalues of the Lindblad map have a non-positive real part, ${\rm Re}\left(\lambda_{k}\right)\leq 0$. Furthermore, trace preservation enforces that at least one eigenvalue is zero, $\lambda_{1}=0$. If such an eigenvalue is non-degenerate—which here we assume to be the case—the (asymptotic) stationary state of the open quantum system,
$\rho_{\rm ss}=\lim_{t\to\infty}\rho(t)\,,$ (64)
@@
Overall, this framework is an exact analog of the “classical final relaxation equilibrium (direct or inverse) Mpemba effect.” To the best of our knowledge, the first manuscript to address the quantum version of the Mpemba effect is [7], where an exactly solvable fully connected quantum Ising model with n-spin exchange with Lindbladian dynamics was considered. Although the framework used in this work is similar to the Lindbladian framework discussed above, an analysis of the effect in terms of the slowest relaxation mode was not given. Instead, it was numerically observed that a (direct) Mpemba effect exist in this model. In [63], a quantum dot system coupled to two thermal baths was theoretically analyzed. A farther analysis of the Mpemba effect in quantum dots that includes electron-electron interactions appears in [101]. A similar framework for the evolution of the density matrix describing the quantum dot was used. Interestingly, it was found that in this specific system $a_{2}=0$ for all initial temperatures because of a symmetry of the problem; hence, the effect is encoded in $a_{3}$ rather than $a_{2}$, and the relevant timescale is $\tau=|\lambda_{3}|^{-1}$, except at the initial condition where $a_{3}=0$ (a strong effect initial condition), where the system relaxes with a characteristic timescale of $\tau=|\lambda_{4}|^{-1}$. A different system that was analyzed with a similar framework and was implemented experimentally in a trapped ions setup was discussed in [8]. In this system, only a strong inverse Mpemba effect was observed, but not a direct effect. The nonequilibrium-steady-state analog of the effect was studied in [102, 62]. In [103, 104], a two-level driven dissipative system subject to an oscillatory electric field and dissipative coupling with the environment [103] and using the trace distance [104] were studied. This system has a transition between real to complex-valued eigenvalues, which turned out to play a role in anomalous relaxation, as it leads to multiple strong Mpemba effects. A perturbative approach to the problem was suggested in [105]. A different approach, based on Davis maps rather than Lindbladiand dynamics, was considered in [106]. The effect was also addressed using information geometry approach in [107]. In [102], the effect was discussed in the famous SYK model, and some connections to quantum chaos were made. A non-Markovian framework was studied in [108], where systems with a finite, non-negligible memory time due to the interactions between the (finite) bath and the system were considered. An analog effect for imaginary time, which is a theoretical tool to evaluate ground-state properties in simulations, was considered in [109]. A similar effect was also discussed in the context of Mosaic models coupled to bosonic environment[110]. A sufficient condition for the existence of the effect was developed in [111]. They introduced a “Mpemba parameter”, which is analogous to $a_{2}(T_{i})$ discussed in the classical case, to quantify how strong the Mpemba effect can be. The case where the initial condition is not a thermal state, but rather some non-equilibrium state, was studied in [112] for the quantum harmonic oscillator coupled to a finite temperature bath. It was demonstrated that any initial distribution of populations with a mean excitation number exactly equal to the mean number of quanta in the thermal equilibrium state displays an accelerated relaxation to equilibrium, namely it corresponds to a strong Mpemba effect. In [113], the issue of how to measure the distance from equilibrium in the quantum setup (in analogy with the discussion in Sec. 2.1.4) was considered. An interesting suggestion to apply the same formalism in photonic setup was given in [114]. In this work, light diffusion in finite-sized photonic lattices under incoherent (dephasing) dynamics is theoretically considered. It is shown that certain highly localized initial light distributions can diffuse faster than initial, broadly delocalized distributions – which is analogous to the classical Mpemba effect in diffusing systems studied, e.g., in [6]. Lastly, in [115], a signature of the Mpemba effect in the context of quantum batteries was identified, where situations of batteries in higher charge states can discharge faster than less charged state was studied.

### v3

@@ 2.1.7 Mpemba effect in boundary and weak coupling setups
Additional analysis that demonstrates the Mpemba effect in this setup can be obtained by projecting the relaxation paths in the (high-dimensional) probability space onto two order parameters, the average $\langle m\rangle$ and staggered magnetization $\langle m_{s}\rangle$ [65]. This projection method, explained in Sec. 7.1.3, allows Monte-Carlo simulation in large systems where a direct calculation of $a_{2}$ is impractical. We can nonetheless compare the relaxation trajectory with the equilibrium distributions at different temperatures; see Sec. 7.1.3 for an example in a 2D Ising antiferromagnet.
Figure 13 illustrates different relaxation paths in a weakly coupled chain of $N=50$ spins in comparison to the equilibrium line. Initial temperatures are chosen before, at, and after the zero of $a_{2}$ at $T_{0}\sim 0.6<T_{b}$. The hot (red) and cold (blue) relaxation paths approach the bath equilibrium at $T_{b}=5$ along the same direction, which is the projection of $\vec{v}_{2}$ into the observable space, but with opposite direction. This corresponds to a sign-change in $a_{2}$, which indeed vanishes at the blue trajectory implying nonmonotonicity of $a_{2}(T_{0})$ (inset of Fig. 13) and consequently a Mpemba effect in the system.
Figure 13: Relaxation paths for a weakly coupled system ($N=50$) on the $m\right>$-$m_{s}\right>$ plane ($H=\sqrt{2}$). Initial temperatures are chosen before (blue), at (green) and after (red) the zero of $a_{2}(T_{0})$ as shown in the inset. Red and green paths approach the equilibrium locus along the same vector but from opposite directions, indicating the existence of a Mpemba effect. Source: Adapted with permission from [70].
@@
Figure 14: (a) Potential $U(x)$ of the colloidal system used in [64]. (b) Examples of Boltzmann distributions at cold and hot bath temperatures for which the direct (cooling) and inverse (heating) Mpemba effect exist. (c) Mpemba phase diagram as a function of the bath temperature $T_{b}$ and thermal diffusivity $\kappa$. Source: Reprinted with permission from [70].
@@
$\partial_{t}p(x,t)=\frac{1}{\gamma}\partial_{x}\Big{[}\big{(}\partial_{x}U(x)\big{)}p(x,t)\Big{]}+\frac{k_{B}}{\gamma}\partial_{x}^{2}\Big{(}T(x,t)p(x,t)\Big{)},$ (47)
where $U(x)$ is the potential and $\gamma$ is the drag coefficient of the Brownian particle. Although the Fokker-Plank operator is time-dependent, the long-time limit $T(x)\to T_{b}$ implies $\mathcal{L}(t)\to\mathcal{L}_{T_{b}}$ and ensures convergence to the Boltzmann equilibrium $\pi(x,T_{b})\propto e^{-U(x)/k_{B}T_{b}}$. The eigenfunctions of $\mathcal{L}_{T_{b}}$, solving $\mathcal{L}_{T_{b}}v_{i}(x,T_{b})=\lambda_{i}v_{i}(x,T_{b})$ with $0=\lambda_{1}>\lambda_{2}\geq\lambda_{3}\geq\dots,$ form a complete basis; as a result,
@@ 2.1.8 Metastability
The maximum amount of work that can be extracted from a nonequilibrium isothermal protocol is the equality case of Eq. (50), which the authors refer to as the extractable work,
@@
At the end of the first stage, the particle has quickly descended into a well. The local equilibrium probability density, $\pi_{{\rm leq}}(x,T,T_{b})$, locally resembles the (global) equilibrium density $\pi(x,{T_{b}})$, but with altered statistical weights in the left and right wells,
@@
The shape of the potential plays an important role in determining whether there is a Mpemba effect. In this discussion, we focused on cases with metastability in the system, but other aspects of the potential’s geometry are also significant. For example, in  [3] the authors propose a one-dimensional potential that exhibits the Mpemba effect. This potential features a metastable state and a ground state with a wide and shallow basin of attraction; see Fig. 6. The energy landscape has both a “fast” and a “slow” direction for relaxation. The relaxation from the metastable state is “sluggish” because of the barrier. Starting at a high temperature is advantageous because most of the statistical weight resides in the broad parts of the potential, allowing for a “fast” relaxation to the ground state.
The presence of metastability significantly affects the type and occurrence of the Mpemba effect, as illustrated in both the spin-glass example in Section 7.1 and the Random Energy Model discussed in Section 5. However, metastability is not a prerequisite for observing the Mpemba effect. Walker and Vucelja in [91] examined an overdamped particle diffusing within a piecewise-continuous potential characterized by energetically flat regions. In these flat regions, the particle can diffuse freely but has a high entropy, making it time-consuming to “escape” from them. This piecewise-continuous potential can be coarse-grained into three states, where the energy levels correspond to different free-energy levels, leading to a form of ”hidden” metastability. A generalization of the above results to piecewise linear potential was done by Biswas, Rajesh, and Pal in [92] as well as in [93]. They investigated analytically a piecewise-linear double potential well [92] (as well as the single well, with and without activity, in [93]) and concluded that neither metastability nor the asymmetry of the potential is necessary or sufficient for the Mpemba effect to occur. A similar conclusion appeared in Kumar’s thesis [94], where it was numerically demonstrated that the Mpemba effect exists in a single-well potential without metastability.
@@
Figure 15: Probability density dynamics in the space of projection coefficients, for the double-well potential described in [6, 75]. (a) Two-dimensional projection of the dynamics onto coefficients $a_{2}$ and $a_{3}$. Red line, labeled $G$, denotes the locus of global equilibrium probability densities (equilibria); green line, labeled $G_{\rm leq}$, denotes the locus of local equilibria. Gray lines denote individual trajectories starting from initial temperatures $T_{0}=\{1000,100,\ldots\}$, labeled by black markers, to the global equilibrium at $T_{b}=1$ (large hollow marker with central dot). The time evolution of the trajectories is indicated by a color gradient from dark to light gray. Time along the local equilibrium locus (green line) is proportional to the relaxation time, as the initial stage is fast. The nonmonotonicity of initial temperatures with respect to length along $G_{\rm leq}$ indicates a Mpemba effect. (b) Three-dimensional projection onto the space of $a_{2}$-$a_{3}$-$a_{4}$ coefficients shows that the red and green curves intersect only at $T_{0}=T_{b}$. Both panels are based on data from experiments [95]. Source: Reprinted with permission from [75].
@@ 2.1.9 The role of dynamics
In some systems, e.g., the 1D antiferromagnetic system discussed above, the existence of some type of Mpemba effect seems natural from the shape of the equilibrium line; see, for example, Fig. 13. It is important to stress, however, that this is not the case: the existence or non-existence of these effects is a consequence of the dynamics and cannot be deduced from the equilibrium properties of the system alone. This is because the effect is based on the projection of the equilibrium distribution on $\vec{v}_{2}$, which is a property of the dynamics, and is not an equilibrium property. In contrast to the equilibrium properties, the exact dynamics are non-universal and affected by many specific issues, such as the precise type of coupling between the system and the bath.
One might imagine that if the equilibrium line is sufficiently complex, the equilibrium projections on $\vec{v}_{2}$ could be nonmonotonic, regardless of the direction of $\vec{v}_{2}$. However, this is never the case: for any system that obeys detailed balance, there is at least one dynamical law in which no Mpemba effect occurs. In fact, a fully connected graph with Metropolis dynamics does not exhibit any type of Mpemba effect, regardless of the equilibrium distribution [96]. Interestingly, in some cases, removing a single link, which corresponds to introducing an infinite barrier, in a fully connected graph with Metropolis dynamics can be enough to induce the Mpemba effect [96]. Similarly, in [97], it was shown that in chemical reactions with identical initial conditions and the same equilibrium target distribution, the existence of the Mpemba effect depends on the precise details of the dynamics. Altering the dynamics, in some cases, one can make the effect appear or disappear. Such implications of dynamics in the context of the rates of chemical reactions are discussed in Sec. 8.1.4. Additionally, applications related to optimal transport and engine efficiency can be found in Sec. 8.1.2.
@@ 2.2 Quantum systems
Until recently, the Mpemba effect was studied exclusively in the context of classical systems. Indeed, a 2018 review of “Quantum quench dynamics” [98] has no mention of a Mpemba or even a Mpemba-related effect. In the same year, though, a phenomena reminiscent of the Mpemba effect was observed in a numerical simulation of a dilute atomic gas in an optical resonator [99]. In that case, there was a faster buildup of the amplitudes of the order-parameter correlation function, for low-frequency modes. Then, in the last five years, the situation changed dramatically, and this review includes over fifty references that relate to the Mpemba effect or analogs seen in quantum systems. A very recent review gives a short overview of these results [100].
It is largely believed that none of the quantum effects are at the origin of the classical effects. Instead, these are quantum effects that share properties that are similar to classical effects. Much as in classical systems, several different quantum effects have been considered, and some were observed experimentally. Below, we give a brief conceptual overview of this fast-developing field. We discuss experiments from 2024 in Sec. 6.
@@ 2.2.1 Theoretical framework: Markovian open quantum systems
Since the dynamics generated by $\mathcal{L}$ is completely positive, all the eigenvalues of the Lindblad map have a non-positive real part, ${\rm Re}\left(\lambda_{k}\right)\leq 0$. Furthermore, trace preservation enforces that at least one eigenvalue is zero, $\lambda_{1}=0$. If such an eigenvalue is non-degenerate, which here we assume to be the case, the (asymptotic) stationary state of the open quantum system
$\rho_{\rm ss}=\lim_{t\to\infty}\rho(t)$ (64)
@@
Overall, this framework is an exact analog of the “classical final relaxation equilibrium (direct or inverse) Mpemba effect.” To the best of our knowledge, the first manuscript to address the quantum version of the Mpemba effect is [7], where an exactly solvable fully connected quantum Ising model with $n$-spin exchange with Lindbladian dynamics was considered. Although the framework used in this work is similar to the Lindbladian framework discussed above, an analysis of the effect in terms of the slowest relaxation mode was not given. Instead, it was numerically observed that a (direct) Mpemba effect exist in this model. In [63], a quantum dot system coupled to two thermal baths was theoretically analyzed. A farther analysis of the Mpemba effect in quantum dots that includes electron-electron interactions appears in [102]. A similar framework for the evolution of the density matrix describing the quantum dot was used. Interestingly, it was found that in this specific system $a_{2}=0$ for all initial temperatures because of a symmetry of the problem; hence, the effect is encoded in $a_{3}$ rather than $a_{2}$, and the relevant timescale is $\tau=|\lambda_{3}|^{-1}$, except at the initial condition where $a_{3}=0$ (a strong effect initial condition), where the system relaxes with a characteristic timescale of $\tau=|\lambda_{4}|^{-1}$. The case of vanishing gap between $\lambda_{2}$ and $\lambda_{3}$ in the same setup was studied in [103], where it was also suggested to accelerate the relaxation by reducing $\lambda_{2}$, through Floquet engineering. A different system that was analyzed with a similar framework and was implemented experimentally in a trapped ions setup was discussed in [8]. In this system, only a strong inverse Mpemba effect was observed, but not a direct effect. The nonequilibrium-steady-state analog of the effect was studied in [104, 62]. In [105, 106], a two-level driven dissipative system subject to an oscillatory electric field and dissipative coupling with the environment [105] and using the trace distance [106] were studied. This system has a transition between real to complex-valued eigenvalues, which turned out to play a role in anomalous relaxation, as it leads to multiple strong Mpemba effects. A perturbative approach to the problem was suggested in [107]. A different approach, investigating the dynamics of Lindbladians with thermal steady states, was considered in [108]. The effect was also addressed using information geometry approach in [109]. In [110], the effect was discussed in the famous Sachdev-Ye-Kitaev (SYK) model, and some connections to quantum chaos were made. A non-Markovian framework was studied in [111], where systems with a finite, non-negligible memory time due to the interactions between the (finite) bath and the system were considered. An analog effect for imaginary time, which is a theoretical tool to evaluate ground-state properties in simulations, was considered in [112]. Additional example of a Mpemba effect in a non-Markovian open quantum system was suggested in [113]. A similar effect was also discussed in the context of Mosaic models coupled to bosonic environment [114]. A sufficient condition for the existence of the effect was developed in [115]. They introduced a “Mpemba parameter”, which is analogous to $a_{2}(T_{i})$ discussed in the classical case, to quantify how strong the Mpemba effect can be. The case where the initial condition is not a thermal state, but rather some nonequilibrium state, was studied in [116] for the quantum harmonic oscillator coupled to a finite temperature bath. It was demonstrated that any initial distribution of populations with a mean excitation number exactly equal to the mean number of quanta in the thermal equilibrium state displays an accelerated relaxation to equilibrium, namely it corresponds to a strong Mpemba effect. In [117], the issue of how to measure the distance from equilibrium in the quantum setup (in analogy with the discussion in Sec. 2.1.4) was considered. An interesting suggestion to apply the same formalism in photonic setup was given in [118, 119]. In the first work [118], light diffusion in finite-sized photonic lattices under incoherent (dephasing) dynamics is theoretically considered. It is shown that certain highly localized initial light distributions can diffuse faster than initial, broadly delocalized distributions. This is analogous to the classical Mpemba effect in diffusing systems studied, e.g., in [6]. In the second work [119], a leaky optical resonator or waveguide was studied, and it was shown that the Mpemba effect can be observed when comparing the decay dynamics of coherent states (classical light) with certain non-classical states. Lastly, in [120], a signature of the Mpemba effect in the context of quantum batteries was identified, where situations of batteries in higher charge states can discharge faster than less charged state was studied.